# Wavelet Analysis of Noninvasive EEG Signals Discriminates Complex and Natural Grasp Types

### v1

@@ I Introduction
Wavelet signal processing has emerged as a powerful tool in analyzing non-stationary signals such as EEGs, and its application in neurophysiology, especially for decoding motor actions, provides a more nuanced understanding of brain activity compared to traditional methods [5]. Specifically, wavelet analysis allows for the extraction of both time and frequency information from EEG data, which is pivotal in identifying neural patterns associated with different types of hand grasps. This dual capability is crucial for identifying specific neural patterns associated with various hand grasps. It addresses a critical gap in existing research, where most studies focus either on time-domain features or frequency-domain features [6, 7, 8], but there are limited studies that consider both of them. To fully interpret EEG signals during complex motor tasks, it’s essential to combine time-domain analysis for timing aspects with frequency-domain analysis for rhythmic activity.
In this study, we applied wavelet signal processing to EEG data to classify complex and natural hand grasp types in a new EEG-based BCI platform. By analyzing the EEG data of participants performing complex power and precision grasps, as well as a no-movement condition, we aimed to identify the unique neural patterns associated with these movements. Machine learning algorithms: Support Vector Machine (SVM), Random Forest (RF), Extreme Gradient Boosting (XGBoost), and Linear Discriminant Analysis (LDA) were utilized to classify these patterns, and permutation feature importance analysis was conducted to determine the most significant features for classification.
@@ II-C Data Recording and Preprocessing
The EEG data in the study were first high-pass filtered at 1Hz to minimize drifts, then bandpass filtered between 1Hz and 30Hz to focus on frequencies relevant to grasping tasks. Artifacts were removed using Independent Component Analysis (ICA), followed by baseline correction using the 200 ms pre-event period. Finally, Z-score normalization was applied to standardize the signals across channels and participants, ensuring data integrity for analysis.
@@ III-B Classification
The classification results in Table 1 show that all four machine learning algorithms achieved accuracies above the 33.33% chance level for three-class classification, indicating the effectiveness of the methods used to classify hand-grasping actions. XGBoost was the most accurate algorithm with an average of 85.16%, closely followed by Random Forest with 83.10%, highlighting the success of these algorithms in interpreting complex EEG data.
TABLE I: 5-fold cross-validated accuracies for multiclass classification
@@
TABLE II: 5-fold cross-validated accuracies for binary classifications
@@
Table 2, encompassing results for various binary classifications, reveals significant findings. In the No-Movement vs Power Grasp classification, all models surpassed the 50% chance level, with SVM leading at 95.37% accuracy and XGBoost closely following. This demonstrates their capability to distinguish between dynamic power grasp and static no-movement conditions. In the No-Movement vs Precision Grasp category, Random Forest emerged as the top performer with an accuracy of 95.40%, with SVM also showing strong results at 94.64%. The classification between Power Grasp and Precision Grasp indicated a general decrease in model accuracies, reflecting the complexity of distinguishing between these grasp types compared to no-movement vs grasp conditions. Here, LDA showed a notable decrease in performance, averaging 70.08% accuracy. However, XGBoost maintained high efficacy, leading with 88.07% accuracy, indicating its robustness in handling complex EEG data classification challenges.
@@ IV Discussion
Finally, the feature importance analysis in Figure 3 shows that key features for grasp classification are primarily in the brain’s motor areas, especially within the alpha and beta bands. These findings align with existing neurophysiological research [15], highlighting the importance of these regions and frequency bands in hand movement. Notably, features such as ’Cz alpha kurtosis’ and ’C3 beta mean’ in central and parietal areas are identified as highly influential.
Overall, the results highlight critical EEG features for grasp type classification and demonstrate the effectiveness of machine learning in interpreting complex neural patterns. The results could potentially be useful for developing high dexterous neuroprosthetic and brain-machine interfaces.

### v2

@@ I Introduction
Wavelet signal processing has emerged as a powerful tool for analyzing non-stationary signals such as EEGs, offering a more nuanced understanding of brain activity in neurophysiology, particularly in decoding motor actions [5]. This method enables the extraction of both time and frequency information from EEG data, which is crucial for identifying neural patterns associated with various types of hand grasps. Our study leverages the capabilities of wavelet analysis to address a critical research gap: while most existing studies concentrate on either time-domain or frequency-domain features [6, 7, 8], there are limited studies that consider the integration of both. By combining these analyses, our research provides a comprehensive interpretation of EEG signals during complex motor tasks, enhancing the classification of different grasp types.
In this study, we utilized wavelet signal processing to analyze EEG data on a new BCI platform, focusing on classifying complex hand grasp types. We extracted novel wavelet-based features to identify unique neural patterns associated with power and precision grasps, as well as a no-movement condition. Using machine learning algorithms such as Support Vector Machine (SVM), Random Forest (RF), Extreme Gradient Boosting (XGBoost), and Linear Discriminant Analysis (LDA), we classified these patterns. A significant contribution of this research is the application of permutation feature importance analysis, which identifies the most critical features and electrode positions.
@@ II-C Data Recording and Preprocessing
The EEG data in the study were bandpass filtered between 1Hz and 30Hz to focus on frequencies relevant to grasping tasks. Artifacts were removed using Independent Component Analysis (ICA), followed by baseline correction using the 200 ms pre-event period. Finally, Z-score normalization was applied to standardize the signals across channels and participants, ensuring data integrity for analysis.
@@ III-B Classification
The classification results in Table 1 show that all four machine learning algorithms achieved accuracies above the 33.33% chance level for three-class classification, indicating the effectiveness of the methods used to classify hand-grasping actions. XGBoost was the most accurate algorithm with an average of 85.16%, closely followed by Random Forest with 83.10%.
TABLE I: Accuracies for multiclass classification
@@
STD 3.51 1.48 2.07 3.06
TABLE II: Accuracies for binary classifications
@@
STD 0.87 2.78 2.36 4.42 1.49 0.59 1.63 2.60 2.34 2.63 4.04 2.67
Table 2, encompassing results for various binary classifications, reveals significant findings. In the No-Movement vs Power Grasp classification, all models surpassed the 50% chance level, with SVM leading at 95.37% accuracy and XGBoost closely following. In the No-Movement vs Precision Grasp category, Random Forest emerged as the top performer with an accuracy of 95.40%, with SVM also showing strong results at 94.64%. The classification between Power Grasp and Precision Grasp indicated a general decrease in model accuracies, reflecting the complexity of distinguishing between these grasp types compared to no-movement vs grasp conditions. Here, LDA showed a notable decrease in performance, averaging 70.08% accuracy. However, XGBoost maintained high efficacy, leading with 88.07% accuracy, indicating its robustness in handling complex EEG data classification challenges.
@@ IV Discussion
Finally, the feature importance analysis in Figure 3 shows that key features for grasp classification are primarily in the brain’s motor areas, especially within the alpha and beta bands. These findings align with existing neurophysiological research [15], highlighting the importance of these regions and frequency bands in hand movement. Notably, features such as ’Cz alpha kurtosis’ and ’C3 beta mean’ in central and parietal areas are identified as highly influential. The ’Cz alpha kurtosis’ reflects transitions between rest and motor execution, crucial for distinguishing between no-movement and active grasping, while the ’C3 beta mean’ captures variations in motor planning and execution, essential for differentiating power from precision grasps.